OpenStudy (anonymous):

Write an equation in slope-intercept form of the line satisfying the given conditions. Through (-7, -9), perpendicular to 4x - 5y = 2

OpenStudy (anonymous):

5x+4y+71=0

OpenStudy (anonymous):

4x - 5y = 2 5y = 4x - 2 y = (4/5) x - 2/5 slope = 4/5 so slope of a line perpendicular to this line = -5/4 the reqd. line also passes thru (-7,-9) using y-y1 = m(x-x1): y- -9 = -(5/4) (x - -7) y+9 = -(5/4)x - 35/4 y = (-5/4)x - 71/4

OpenStudy (anonymous):

y=(-5/4)x+(-71/4)

OpenStudy (anonymous):

-71/4 should be in brakets